OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

please

OpenStudy (luigi0210):

Solve for x first then plug that in

OpenStudy (anonymous):

You could use substitution or elimination.

OpenStudy (anonymous):

is this the answer (-3, 1)

OpenStudy (nurali):

x - 4y = 1----(1) 2x - 9y = 3---(2) in equation (1) x=1+4y put in equation (2) 2(1+4y)-9y=3 2+8y-9y=3 2-y=3 y=-1 put in equation (1) x-4(-1)=1 x+4=1 x=-3 so the solution is (x,y)=(-3,-1)

OpenStudy (wolf1728):

Sorry messed up a bit there: A) x - 4y = 1 B) 2x - 9y = 3 Solve equation A for x x = 1 + 4y then substitute this into Equation B